# A. I. Bolotnikov, A. A. Irmatov

### v1

@@ Introduction
The value for $k_{1}$ can be chosen in $p-1$ ways. Let us consider the following cases:
  • Case 1: $k_{3}=k_{1}$. The conditions for choosing a value for $k_{2}$ are $k_{2}\neq 0,k_{2}\neq k_{1},k_{2}\neq 2k_{1}$, and we can choose $k-2$ in $p-3$ ways. In this case there are 4 distinct values among $0,k_{1},k_{2},k_{3},k_{3}+k_{1}-k_{2}$ therefore there are $(p-1)(p-3)(p-4)...(p-n)$ ways to choose values for $(x_{1},x_{2},...,x_{n})$ in this case.
  • Case 2: $k_{2}=2k_{1},k_{3}=3k_{1}$. There are 4 distinct values among $0,k_{1},k_{2},k_{3},\\ k_{3}+k_{1}-k_{2}$ therefore there are $(p-1)(p-4)(p-5)...(p-n)$ ways to choose values for $(x_{1},x_{2},...,x_{n})$ in this case.
  • Case 3: $k_{2}=2k_{1},k_{3}\neq 3k_{1},k_{3}\neq k_{1}$. The conditions for choosing a value for $k_{3}$ are $k_{3}\neq 0,k_{3}\neq k_{1},k_{3}\neq k_{2},k_{3}\neq 3k_{1}$, and we can choose $k_{3}$ in $p-4$ ways. There are 5 distinct values among $0,k_{1},k_{2},k_{3},k_{3}+k_{1}-k_{2}$ , therefore there are $(p-1)(p-4)(p-5)(p-6)...(p-(n+1))$ ways to choose values for $(x_{1},x_{2},...,x_{n})$ in this case .
  • Case 4: $k_{2}=k_{1}/2,k_{3}\neq k_{1}$. The conditions for choosing a value for $k_{3}$ are $k_{3}\neq 0,k_{3}\neq k_{1},k_{3}\neq k_{2},k_{3}\neq k_{2}-k_{1}$, and we can choose $k_{3}$ in $p-4$ ways. There are 5 distint values among $0,k_{1},k_{2},k_{3},k_{3}+k_{1}-k_{2}$ therefore there are $(p-1)(p-4)(p-5)(p-6)...(p-(n+1))$ to choose values for $(x_{1},x_{2},...,x_{n})$ in this case.
  • Case 5: $k_{2}\neq 2k_{1},k_{2}\neq k_{1}/2,k_{3}=2k_{2}-k_{1}$. The conditions for choosing a value for $k_{2}$ are $k_{2}\neq 0,k_{2}\neq k_{1},k_{2}\neq 2k_{1},k_{2}\neq k_{1}/2$, and we can choose $k_{2}$ in $p-4$ ways.There are 4 distinct values among $0,k_{1},k_{2},k_{3},k_{3}+k_{1}-k_{2}$, therefore there are $(p-1)(p-4)(p-4)(p-5)...(p-n)$ ways to choose values for $(x_{1},x_{2},...,x_{n})$ in this case.
  • Case 6: $k_{2}\neq 2k_{1},k_{2}\neq k_{1}/2,k_{3}\neq 2k_{2}-k_{1},k_{3}\neq k_{1}$. The conditions for choosing a value for $k_{2}$ are $k_{2}\neq 0,k_{2}\neq k_{1},k_{2}\neq 2k_{1},k_{2}\neq k_{1}/2$, and we can choose $k_{2}$ in $p-4$ ways. The conditions for choosing a value for $k_{3}$ are $k_{3}\neq 0,k_{3}\neq k_{1},k_{3}\neq k_{2},k_{3}\neq k_{2}-k_{1},k_{3}\neq 2k_{2}-k_{1}$ , and we can choose $k_{3}$ in $p-5$ ways. There are 5 distinct values among $0,k_{1},k_{2},k_{3},k_{3}+k_{1}-k_{2}$, therefore there are $(p-1)(p-4)(p-5)(p-5)(p-6)...(p-(n+1))$ ways to choose values for $(x_{1},x_{2},...,x_{n})$ in this case.
The value of $\chi_{MA(G)}(p)$ is the sum of numbers from these cases. As a result, if $G$ consists of a $K_{3}$ subgraph and a “tail” subgraph that are connected through a shared vertex, the formula for $\chi_{MA(G)}(x)$ is $\chi_{MA(G)}(x)=(x-1)(x-3)(x-4)....(x-(n-2))(x-(n-1))(x-(n-1))(x-n)$. ∎
@@
Let $G(V,E),|E|=n$ be a graph without loops and parallel edges. Let $F_{p}$ be the group of residue classes modulo a prime number. A weight function $W:E\rightarrow F_{p}^{k}$ is called improper, if there is a sequence of edges $(e_{i_{1}},e_{i_{2}},...,e_{i_{d}})$ that form a simple path or a simple even cycle in G, such that
$W(e_{i_{1}})-W(e_{i_{2}})+W(e_{i_{3}})-....+(-1)^{d-1}W(e_{i_{d}})={\bf 0}\in F_{p}^{k}.$ (1)
Otherwise the weight function is called proper. The choice of names is motivated by the following consideration. If a weight function on the edges of graph $G$ is given, then the weight of a matching can be defined as the sum of the weights of its edges. Furthermore, there is a one-to-one correspondence between the matchings of graph G and the vertices of the matching polytope of graph G [8]. Thus, based on the weight function on the edges of graph G, it is possible to construct a weight function on the vertices of the matching polytope of graph G. An important property of the matching polytope is that any two vertices of this polytope are connected with the edge, if an only if the symmetrical difference of the corresponding matchings consists of exactly one simple path of arbitrary length or one simple cycle of even length [9, Theorem 25.3]. If the original weight function on the edges of the graph was proper, then for the resulting weight function on the vertices of the polytope, any two vertices connected by an edge of the polytope have different weights. Conversely, in case of an improper initial weight function on the edges of the graph, at least one pair of vertices with the same weight connected by an edge will be found in the polytope. In this case the path or the cycle that satisfies (1) is the symmetrical difference of two matchings that correspond to vertices of the same weight in the matching polytope that are connected by an edge.
Let us state the improper weight function problem. For an arbitrary graph $G(V,E)$ without loops and parallel edges, an arbitrary prime number p, an arbitrary natural number k and an arbitrary weight function $W:E\rightarrow F_{p}^{k}$ the problem is to establish, if function W is an improper one.
@@
We will first prove that the problem belongs to the complexity class NP. By definition, a recognition problem $A$ with possible answers “Yes” or “No”  belongs to the class NP if, when $A$ has an answer “Yes”, there exists a “hint” $c(A)$, the length of which polynomially depends on the size of $A$, such that the task $(c(A),A)$ is solved in polynomial time.
For a improper weight function problem, let us choose a simple path or a simple even cycle $(e_{i_{1}},e_{i_{2}},...,e_{i_{d}})$ that satisfies (1), as a “hint”. With such a hint the solution of a improper weight function problem is reduced to computing $W(e_{i_{1}})-W(e_{i_{2}})+W(e_{i_{3}})-....+(-1)^{d-1}W(e_{i_{d}})$. Since there are no more then n edges in this path, the computation time depends polynomially from n and k.
To show that improper weight function problem is an NP-complete one, we now transfrom an arbitrary 3-satisfiability problem into an improper weight function problem. Let $X=\{x_{1},x_{2},...,x_{n}\}$ be a set of boolean variables and let $U=\{u_{1},u_{2},...,u_{m}\}$ be set of clauses making up an arbitrary instance of 3-satisfiability problem. The graph G for the improper weight function problem will be constructed in the following way. For each variable $x_{i}$, there will be a subgraph $Z_{i}$ of the graph G that consists of one cycle of length 4, with the set of edges $\{a_{i},b_{i},c_{i},d_{i}\}$ and the set of vertices $\{A_{i},B_{i},C_{i},D_{i}\}$, such that $a_{i}=(A_{i},B_{i}),b_{i}=(B_{i},C_{i}),c_{i}=(C_{i},D_{i}),d_{i}=(D_{i},A_{i})$.
@@
To construct the graph G, we connect $Z_{i}$ with $Z_{i+1}$ through a shared vertex, starting with $Z_{1}$: $C_{i}=A_{i+1}$. Then we connect $Z_{n}$ with $Y_{1}$ through a shared vertex: $C_{n}=P_{1}$. Then we connect $Y_{j}$ with $Y_{j+1}$ through a shared vertex, starting with $Y_{1}$: $Q_{i}=P_{i+1}$.
@@
The constructed graph $G(V,E)$ consists of $13n+8m+$ 1 vertices and $4n+14m$ edges. The next step is to construct the weight function $W:E\rightarrow F_{p}^{k}$. The prime number p can be chosen aribtrarily. Let $k=3n+11m-1$. The values of the weight function $W(e)=(w_{1}(e),....,w_{k}(e))$ will be defined in the following way.
@@
Let $\hat{V}=V\setminus\{A_{1},Q_{m}\}$. $|\hat{V}|=3n+8m-1$. Let all the elements of $\hat{V}$ be numbered: $\hat{V}=\{v_{1},...,v_{3n+8m-1}\}$. For $e\in E$ and $11\leq i\leq 3n+8m-$ 1, let $w_{i}(e)=1$, if $e$ is an adjacent edge to $v_{i}$, and let $w_{i}(e)=0$ otherwise.
Let $u_{j}=(t_{\alpha},t_{\beta},t_{\gamma})$ be a clause.
$w_{3n+8m-1+3\cdot(j-1)+1}(e)=1$, if either $t_{\alpha}=x_{\alpha}$ and $e=(A_{\alpha},B_{\alpha})$, or $t_{\alpha}=\neg x_{\alpha}$ and $e=(A_{\alpha},D_{\alpha})$. $w_{3n+8m-1+3\cdot(j-1)+1}(e)=-1$, if $e\in\{(Q_{j},R_{j,1}),(Q_{j},R_{j,2}),\\ (Q_{j},R_{j,3}),(Q_{j},R_{j,4})\}$. Otherwise, $w_{3n+8m-1+3\cdot(j-1)+1}(e)=0$.
$w_{3n+8m-1+3\cdot(j-1)+2}(e)=1$, if either $t_{\beta}=x_{\beta}$ and $e=(A_{\beta},B_{\beta})$, or $t_{\beta}=\neg x_{\beta}$ and $e=(A_{\beta},D_{\beta})$. $w_{3n+8m-1+3\cdot(j-1)+2}(e)=-1$, if $e\in\{(Q_{j},R_{j,1}),(Q_{j},R_{j,2}),\\ (Q_{j},R_{j,5}),(Q_{j},R_{j,6})\}$. Otherwise, $w_{3n+8m-1+3\cdot(j-1)+2}(e)=0$.
$w_{3n+8m-1+3\cdot(j-1)+3}(e)=1$, if either $t_{\gamma}=x_{\gamma}$ and $e=(A_{\gamma},B_{\gamma})$, or $t_{\gamma}=\neg x_{\gamma}$ and $e=(A_{\gamma},D_{\gamma})$. $w_{3n+8m-1+3\cdot(j-1)+3}(e)=-1$, if $e\in\{(Q_{j},R_{j,1}),(Q_{j},R_{j,3}),\\ (Q_{j},R_{j,5}),(Q_{j},R_{j,7})\}$. Otherwise, $w_{3n+8m-1+3\cdot(j-1)+3}(e)=0$.
@@
Now function W is completely defined. Because of the way that $w_{i}(e),1\leq i\leq 3n+8-1$, were defined, if a path start or ends in any vertex $v_{j}$ from $V\setminus\{A_{1},Q_{m}\}$, there is exactly one edge $e$ on this path that has $w_{j}(e)\neq 0$. Therefore, it is impossible for a simple path $(e_{i_{1}},e_{i_{2}},...,e_{i_{d}})$ that satisfies (1) to have any ends, other then $A_{1}$ and $Q_{m}$. Note that the only cycles of even length in $G$ are cycles of length for inside $Z_{i}$ or $Y_{j}$. It is not possible for a cycle of length 4 inside $Z_{i}$ or $Y_{j}$ to satisfy (1) as well, because with these values for $w_{i}(e),i>3n+8-1$, it is nessesary for every path or even cycle $(e_{i_{1}},e_{i_{2}},...,e_{i_{d}})$ that satisfies (1) and contains an edge in one of $Z_{i}$ to also contain an edge in one of $Y_{j}$, and visa versa. In case of $Z_{i}$ the reason for this is that there is a specific index $t$ that corresponds to a pair $(x_{i},u)$, where $u$ is a clause that contains either $x_{i}$ or $\neg x_{i}$, such that there is exactly one edge $e$ in $Z_{i}$ with $w_{t}(e)\neq 0$. In case of $Y_{j}$, each of the edges $e\in\{(Q_{j},R_{j,1}),(Q_{j},R_{j,2}),(Q_{j},R_{j,3}),(Q_{j},R_{j,4}),(Q_{j},R_{j,5}),(Q_{j},R_{j,6}),(Q_{j},R_{j,7})\}$ has its own distinct vector of values for $(w_{3n+8m-1+3\cdot(j-1)+1}(e),w_{3n+8m-1+3\cdot(j-1)+2}(e),\\ w_{3n+8m-1+3\cdot(j-1)+3}(e))$, and since any cycle of length 4 in $Y_{j}$ contains exatly two of those edges, it can’t satisfy (1). Therefore, the only paths or even cycles that can possibly satisfy (1) are paths that connect $A_{1}$ and $Q_{m}$. Note that k, the size of G and the computation of W are all polynomial from m and n.
@@
Let $\{\alpha_{1},\alpha_{2},...,\alpha_{n}\}$ be a set of values assigned to $X$, such that all clauses $u_{1},u_{2},...,u_{n}$ are satisfied. The current goal is to find a path $(e_{i_{1}},e_{i_{2}},...,e_{i_{d}})$ in G that satisfies (1) knowing $\{\alpha_{1},\alpha_{2},...,\alpha_{n}\}$. Let us consider the following path $(e_{i_{1}},e_{i_{2}},...,e_{i_{2\cdot(m+n)}})$ of length $2\cdot(m+n)$ from $A_{1}$ to $Q_{m}$. For $1\leq i\leq n$, if $\alpha_{i}=True$, then $e_{2\cdot i-1}=a_{i},e_{2\cdot i}=b_{i}$, and if $\alpha_{i}=False$, then $e_{2\cdot i-1}=d_{i},e_{2\cdot i}=c_{i}$.
@@
Let $u_{j}=(t_{\alpha},t_{\beta},t_{\gamma})$ be a clause. Let’s consider all possible cases, after values $\alpha_{1},\alpha_{2},...,\alpha_{n}$ are assigned.
@@
Now the path $P=(e_{i_{1}},e_{i_{2}},...,e_{i_{2\cdot(m+n)}})$ is completely defined. Let $W(e_{i_{1}})-W(e_{i_{2}})+W(e_{i_{3}})-....+(-1)^{2\cdot(m+n)-1}W(e_{i_{2\cdot(m+n)}})=(z_{1},z_{2},...,z_{k})\in F_{p}^{k}$. For $11\leq i\leq 3n+8m-$ 1, $z_{i}=0$, because for every vertex $v\in V\setminus\{A_{1},Q_{m}\}$ path P either passes through $v$ without stopping there or doesn’t pass through $v$ at all. For $i>3n+8m-1$, $z_{i}$ corresponds to a pair $(x_{q},u_{p})$, where $u_{p}$ contains either $x_{q}$ or $\neg x_{q}$. There are four possible cases:
@@
  • Case 1: $\alpha_{q}=True$, $u_{p}$ contains $x_{q}$. In this case $w_{i}(e_{i_{2\cdot q-1}})=1,w_{i}(e_{i_{2\cdot n+2\cdot p-1}})=-1,w_{i}(e)=0$ for every other edge e on P.
@@
  • Case 2: $\alpha_{q}=False$, $u_{p}$ contains $x_{q}$. In this case $w_{i}(e_{i_{2\cdot q-1}})=0,w_{i}(e_{i_{2\cdot n+2\cdot p-1}})=0,w_{i}(e)=0$ for every other edge e on P.
@@
  • Case 3: $\alpha_{q}=True$, $u_{p}$ contains $\neg x_{q}$. In this case $w_{i}(e_{i_{2\cdot q-1}})=0,w_{i}(e_{i_{2\cdot n+2\cdot p-1}})=0,w_{i}(e)=0$ for every other edge e on P.
@@
  • Case 4: $\alpha_{q}=False$, $u_{p}$ contains $\neg x_{q}$. In this case $w_{i}(e_{i_{2\cdot q-1}})=1,w_{i}(e_{i_{2\cdot n+2\cdot p-1}})=-1,w_{i}(e)=0$ for every other edge e on P.
@@
In each of these cases, $z_{i}=0$. Therefore, for a path $P=(e_{i_{1}},e_{i_{2}},...,e_{i_{2\cdot(m+n)}})$, where $e_{i_{1}}$ is adjacent to $A_{1}$ and $e_{i_{2\cdot(m+n)}}$ is adjacent to $Q_{m}$, we have $W(e_{i_{1}})-W(e_{i_{2}})+W(e_{i_{3}})-....+(-1)^{2\cdot(m+n)-1}W(e_{i_{2\cdot(m+n)}})=0$.

### v3

@@ Introduction
The value for $k_{1}$ can be chosen in $p-1$ ways. Let us consider the following cases.
  • Case 1: $k_{3}=k_{1}$. The conditions for choosing a value for $k_{2}$ are $k_{2}\neq 0,k_{2}\neq k_{1},k_{2}\neq 2k_{1}$, and we can choose $k_{2}$ in $p-3$ ways. In this case, there are 4 distinct values among $0,k_{1},k_{2},k_{3},k_{3}+k_{1}-k_{2}$, therefore, there are $(p-1)(p-3)(p-4)...(p-n)$ ways to choose values for $(x_{1},...,x_{n})$ in this case.
  • Case 2: $k_{2}=2k_{1},k_{3}=3k_{1}$. There are 4 distinct values among $0,k_{1},k_{2},k_{3},k_{3}+k_{1}-k_{2},$ therefore, there are $(p-1)(p-4)(p-5)...(p-n)$ ways to choose values for $(x_{1},...,x_{n})$ in this case.
  • Case 3: $k_{2}=2k_{1},k_{3}\neq 3k_{1},k_{3}\neq k_{1}$. The conditions for choosing a value for $k_{3}$ are $k_{3}\neq 0,k_{3}\neq k_{1},k_{3}\neq k_{2},k_{3}\neq 3k_{1}$, and we can choose $k_{3}$ in $p-4$ ways. There are 5 distinct values among $0,k_{1},k_{2},k_{3},k_{3}+k_{1}-k_{2}$, therefore, there are $(p-1)(p-4)(p-5)(p-6)...(p-(n+1))$ ways to choose values for $(x_{1},...,x_{n})$ in this case.
  • Case 4: $k_{2}=k_{1}/2,k_{3}\neq k_{1}$. The conditions for choosing a value for $k_{3}$ are $k_{3}\neq 0,k_{3}\neq k_{1},k_{3}\neq k_{2},k_{3}\neq k_{2}-k_{1}$, and we can choose $k_{3}$ in $p-4$ ways. There are 5 distinct values among $0,k_{1},k_{2},k_{3},k_{3}+k_{1}-k_{2}$, therefore, there are $(p-1)(p-4)(p-5)(p-6)...(p-(n+1))$ ways to choose values for $(x_{1},...,x_{n})$ in this case.
  • Case 5: $k_{2}\neq 2k_{1},k_{2}\neq k_{1}/2,k_{3}=2k_{2}-k_{1}$. The conditions for choosing a value for $k_{2}$ are $k_{2}\neq 0,k_{2}\neq k_{1},k_{2}\neq 2k_{1},k_{2}\neq k_{1}/2$, and we can choose $k_{2}$ in $p-4$ ways. There are 4 distinct values among $0,k_{1},k_{2},k_{3},k_{3}+k_{1}-k_{2}$, therefore, there are $(p-1)(p-4)(p-4)(p-5)...(p-n)$ ways to choose values for $(x_{1},...,x_{n})$ in this case.
  • Case 6: $k_{2}\neq 2k_{1},k_{2}\neq k_{1}/2,k_{3}\neq 2k_{2}-k_{1},k_{3}\neq k_{1}$. The conditions for choosing a value for $k_{2}$ are $k_{2}\neq 0,k_{2}\neq k_{1},k_{2}\neq 2k_{1},k_{2}\neq k_{1}/2$, and we can choose $k_{2}$ in $p-4$ ways. The conditions for choosing a value for $k_{3}$ are $k_{3}\neq 0,k_{3}\neq k_{1},k_{3}\neq k_{2},k_{3}\neq k_{2}-k_{1},k_{3}\neq 2k_{2}-k_{1}$, and we can choose $k_{3}$ in $p-5$ ways. There are 5 distinct values among $0,k_{1},k_{2},k_{3},k_{3}+k_{1}-k_{2}$, therefore, there are $(p-1)(p-4)(p-5)(p-5)(p-6)...(p-(n+1))$ ways to choose values for $(x_{1},...,x_{n})$ in this case.
The value of $\chi_{MA(G)}(p)$ is the sum of numbers from these cases. As a result, if $G$ consists of a $K_{3}$ subgraph and a “tail” subgraph that are connected through a shared vertex, the formula for $\chi_{MA(G)}(x)$ is
$\chi_{MA(G)}(x)=(x-1)(x-3)(x-4)...(x-(n-2))(x-(n-1))(x-(n-1))(x-n).$
@@
Let $G(V,E),$ $|E|=n$ be a graph without loops, multiple edges and isolated vertices. Let $F_{p}$ be the group of residue classes modulo a prime number $p$. A weight function $W:E\rightarrow F_{p}^{k}$ with an arbitrary $k\in\mathbb{N}$ is called improper, if there is a sequence of edges $(e_{i_{1}},...,e_{i_{d}})$ that form a simple path or a simple even cycle in $G$, such that
$W(e_{i_{1}})-W(e_{i_{2}})+W(e_{i_{3}})-...+(-1)^{d-1}W(e_{i_{d}})={\bf 0}\in F_{p}^{k}.$ (1)
Otherwise the weight function is called proper. The choice of names is motivated by the following consideration. If a weight function on the edges of graph $G$ is given, then the weight of a matching can be defined as the sum of the weights of its edges. Furthermore, there is a one-to-one correspondence between the matchings of graph $G$ and the vertices of the matching polytope of graph $G$ [8]. Thus, based on the weight function on the edges of graph $G$, it is possible to construct a weight function on the vertices of the matching polytope of graph $G$. An important property of the matching polytope is that any two vertices of this polytope are connected with the edge, if and only if the symmetrical difference of the corresponding matchings consists of exactly one simple path of arbitrary length or one simple cycle of even length [9, Theorem 25.3].
If the original weight function on the edges of the graph was proper, then for the resulting weight function on the vertices of the polytope, any two vertices connected by an edge of the polytope have different weights. Conversely, if the initial weight function on the edges of the graph was improper, there is at least one pair of vertices in the matching polytope that are connected by an edge and have the same weight. In this case, the path or the cycle that satisfies (1) is the symmetrical difference of two matchings that correspond to these two vertices.
Let us state the improper weight function problem. For an arbitrary graph $G(V,E),|E|=n,$ without loops, multiple edges and isolated vertices, an arbitrary prime number $p$, an arbitrary natural number $k$ and an arbitrary weight function $W:E\rightarrow F_{p}^{k}$ the problem is to establish, if function $W$ is an improper one.
@@
We will first prove that the problem belongs to the complexity class NP. By definition, a recognition problem $A$ with possible answers “Yes” or “No” belongs to the class NP if there is a nondeterministic algorithm that consists of a guessing stage and a checking stage, such that when $A$ has an answer “Yes”, the guessing stage returns a structure $S$, whose size depends polynomially from the size of the input to $A$, and then the checking stage takes $S$ and the input to $A$ as its own input and verifies in polynomial time, that the answer to $A$ was “Yes”.
For an improper weight function problem, an input consists of graph $G$ that contains $n$ edges and $b,b\in\mathbb{N},b(b-1)/2\leq n$ vertices, and a weight function $W:E\rightarrow F_{p}^{k}$ that can be described with $nk$ elements of $F_{p}$. The size of an input to an improper weight function problem is $O(nk)$. Given an instance of an improper weight function problem with answer “Yes”, a nondetermininstic algorithm can guess the path $(e_{i_{1}},...,e_{i_{d}})$ that satisfies (1), and then compute $W(e_{i_{1}})-W(e_{i_{2}})+W(e_{i_{3}})-...+(-1)^{d-1}W(e_{i_{d}})$ to verify, that the answer to the improper weight function problem is “Yes”. And since $d\leq n$, both the size of structure returned on the guessing stage and the time of computation on the checking stage are polynomial from the input to the imporper weight function problem.
To show that improper weight function problem is an NP-complete one, we now transform an arbitrary 3-satisfiability problem into an improper weight function problem. Let $X=\{x_{1},...,x_{n}\}$ be a set of Boolean variables and let $U=\{u_{1},...,u_{m}\}$ be set of clauses making up an arbitrary instance of 3-satisfiability problem. The graph $G$ for the improper weight function problem will be constructed in the following way. For each variable $x_{i}$, there will be a subgraph $Z_{i}$ of the graph $G$ that consists of one cycle of length 4, with the set of edges $\{a_{i},b_{i},c_{i},d_{i}\}$ and the set of vertices $\{A_{i},B_{i},C_{i},D_{i}\}$, such that $a_{i}=(A_{i},B_{i}),b_{i}=(B_{i},C_{i}),c_{i}=(C_{i},D_{i}),d_{i}=(D_{i},A_{i})$.
@@
To construct the graph $G$, we connect $Z_{i}$ with $Z_{i+1}$ through a shared vertex, starting with $Z_{1}$: $C_{i}=A_{i+1}$. Then, we connect $Z_{n}$ with $Y_{1}$ through a shared vertex: $C_{n}=P_{1}$. Then, we connect $Y_{j}$ with $Y_{j+1}$ through a shared vertex, starting with $Y_{1}$: $Q_{i}=P_{i+1}$.
@@
The constructed graph $G(V,E)$ consists of $13n+8m+$ 1 vertices and $4n+14m$ edges. The next step is to construct the weight function $W:E\rightarrow F_{p}^{k}$. The prime number $p$ can be chosen arbitrarily. Let $k=3n+11m-1$. The values of the weight function $W(e)=(w_{1}(e),...,w_{k}(e))$ will be defined in the following way.
@@
Let $\hat{V}=V\setminus\{A_{1},Q_{m}\}$. $|\hat{V}|=3n+8m-1$. Let all the elements of $\hat{V}$ be numbered: $\hat{V}=\{v_{1},...,v_{3n+8m-1}\}$. For $e\in E$ and $11\leq i\leq 3n+8m-$ 1, let $w_{i}(e)=1$, if $e$ is an adjacent edge to $v_{i}$, and let $w_{i}(e)=0$ otherwise. Let $u_{j}=(t_{\alpha},t_{\beta},t_{\gamma})$ be a clause.
$w_{3n+8m-1+3(j-1)+1}(e)=1$, if either $t_{\alpha}=x_{\alpha}$ and $e=(A_{\alpha},B_{\alpha})$, or $t_{\alpha}=\neg x_{\alpha}$ and $e=(A_{\alpha},D_{\alpha})$.
$w_{3n+8m-1+3(j-1)+1}(e)=-1$, if $e\in\{(Q_{j},R_{j,1}),(Q_{j},R_{j,2}),(Q_{j},R_{j,3}),(Q_{j},R_{j,4})\}$. Otherwise, $w_{3n+8m-1+3(j-1)+1}(e)=0$.
$w_{3n+8m-1+3(j-1)+2}(e)=1$, if either $t_{\beta}=x_{\beta}$ and $e=(A_{\beta},B_{\beta})$, or $t_{\beta}=\neg x_{\beta}$ and $e=(A_{\beta},D_{\beta})$.
$w_{3n+8m-1+3(j-1)+2}(e)=-1$, if $e\in\{(Q_{j},R_{j,1}),(Q_{j},R_{j,2}),(Q_{j},R_{j,5}),(Q_{j},R_{j,6})\}$. Otherwise, $w_{3n+8m-1+3(j-1)+2}(e)=0$.
$w_{3n+8m-1+3(j-1)+3}(e)=1$, if either $t_{\gamma}=x_{\gamma}$ and $e=(A_{\gamma},B_{\gamma})$, or $t_{\gamma}=\neg x_{\gamma}$ and $e=(A_{\gamma},D_{\gamma})$.
$w_{3n+8m-1+3(j-1)+3}(e)=-1$, if $e\in\{(Q_{j},R_{j,1}),(Q_{j},R_{j,3}),(Q_{j},R_{j,5}),(Q_{j},R_{j,7})\}$. Otherwise, $w_{3n+8m-1+3(j-1)+3}(e)=0$.
@@
Now function $W$ is completely defined. Because of the way that $w_{i}(e),1\leq i\leq 3n+8-1$, were defined, if a path starts or ends in any vertex $v_{j}$ from $V\setminus\{A_{1},Q_{m}\}$, there is exactly one edge $e$ on this path that has $w_{j}(e)\neq 0$. Therefore, it is impossible for a simple path $(e_{i_{1}},...,e_{i_{d}})$ that satisfies (1) to have any ends, other then $A_{1}$ and $Q_{m}$. Note that the only cycles of even length in $G$ are cycles of length 4 inside $Z_{i}$ or $Y_{j}$. It is not possible for a cycle of length 4 inside $Z_{i}$ or $Y_{j}$ to satisfy (1) as well, because with these values for $w_{i}(e),i>3n+8-1$, it is nessesary for every path or even cycle $(e_{i_{1}},...,e_{i_{d}})$ that satisfies (1) and contains an edge in one of $Z_{i}$ to also contain an edge in one of $Y_{j}$, and visa versa. In case of $Z_{i}$, the reason for this is that there is a specific index $t$ that corresponds to a pair $(x_{i},u)$, where $u$ is a clause that contains either $x_{i}$ or $\neg x_{i}$, such that there is exactly one edge $e$ in $Z_{i}$ with $w_{t}(e)\neq 0$. In case of $Y_{j}$, each of the edges $e\in\{(Q_{j},R_{j,1}),(Q_{j},R_{j,2}),(Q_{j},R_{j,3}),(Q_{j},R_{j,4}),(Q_{j},R_{j,5}),(Q_{j},R_{j,6}),(Q_{j},R_{j,7})\}$ has its own distinct vector of values for $(w_{3n+8m-1+3(j-1)+1}(e),w_{3n+8m-1+3(j-1)+2}(e),w_{3n+8m-1+3(j-1)+3}(e))$, and since any cycle of length 4 in $Y_{j}$ contains exactly two of those edges, it can’t satisfy (1). Therefore, the only paths or even cycles that can possibly satisfy (1) are paths that connect $A_{1}$ and $Q_{m}$. Note that $k$, the size of $G$ and the computation of $W$ are all polynomial from $m$ and $n$.
@@
Let $\{\alpha_{1},...,\alpha_{n}\}$ be a set of values assigned to $X$, such that all clauses $u_{1},...,u_{n}$ are satisfied. The current goal is to find a path $(e_{i_{1}},...,e_{i_{d}})$ in $G$ that satisfies (1), knowing $\{\alpha_{1},...,\alpha_{n}\}$. Let us consider the following path $(e_{i_{1}},...,e_{i_{2(m+n)}})$ of length $2(m+n)$ from $A_{1}$ to $Q_{m}$. For $1\leq i\leq n$, if $\alpha_{i}=True$, then $e_{2i-1}=a_{i},e_{2i}=b_{i}$, and if $\alpha_{i}=False$, then $e_{2i-1}=d_{i},e_{2i}=c_{i}$.
@@
Let $u_{j}=(t_{\alpha},t_{\beta},t_{\gamma})$ be a clause. Let us consider all possible cases, after values $\alpha_{1},...,\alpha_{n}$ are assigned.
@@
For $11\leq i\leq 3n+8m-$ 1, $z_{i}=0$, because for every vertex $v\in V\setminus\{A_{1},Q_{m}\}$ path $P$ either passes through $v$ without stopping there or doesn’t pass through $v$ at all. For $i>3n+8m-1$, $z_{i}$ corresponds to a pair $(x_{q},u_{p})$, where $u_{p}$ contains either $x_{q}$ or $\neg x_{q}$. There are four possible cases:
@@
  • Case 1: $\alpha_{q}=True$, $u_{p}$ contains $x_{q}$. In this case, $w_{i}(e_{i_{2q-1}})=1,w_{i}(e_{i_{2n+2p-1}})=-1,w_{i}(e)=0$ for every other edge $e$ on $P$.
@@
  • Case 2: $\alpha_{q}=False$, $u_{p}$ contains $x_{q}$. In this case, $w_{i}(e_{i_{2q-1}})=0,w_{i}(e_{i_{2n+2p-1}})=0,w_{i}(e)=0$ for every other edge $e$ on $P$.
@@
  • Case 3: $\alpha_{q}=True$, $u_{p}$ contains $\neg x_{q}$. In this case, $w_{i}(e_{i_{2q-1}})=0,w_{i}(e_{i_{2n+2p-1}})=0,w_{i}(e)=0$ for every other edge $e$ on $P$.
@@
  • Case 4: $\alpha_{q}=False$, $u_{p}$ contains $\neg x_{q}$. In this case, $w_{i}(e_{i_{2q-1}})=1,w_{i}(e_{i_{2n+2p-1}})=-1,w_{i}(e)=0$ for every other edge $e$ on $P$.
@@
In each of these cases, $z_{i}=0$. Therefore, for a path $P=(e_{i_{1}},...,e_{i_{2(m+n)}})$, where $e_{i_{1}}$ is adjacent to $A_{1}$ and $e_{i_{2(m+n)}}$ is adjacent to $Q_{m}$, we have
$W(e_{i_{1}})-W(e_{i_{2}})+W(e_{i_{3}})-...+(-1)^{2(m+n)-1}W(e_{i_{2(m+n)}})=0.$
Conversely, let $P=(e_{i_{1}},...,e_{i_{2(m+n)}})$ be a path, such that
$W(e_{i_{1}})-W(e_{i_{2}})+W(e_{i_{3}})-...+(-1)^{2(m+n)-1}W(e_{i_{2(m+n)}})=0.$